Prove that the set W = {(x, y, z) * + = 0} is a subspace of Rs and then find a basis in W.
Answer #1

8 a W = { (3,4,7) | 2+z=0} 0+0=0 (0,0,0) EW 020, E W, le ß : (x2, 9₂, Z) Let & (x,,Y,Z) e & a ER. then, 2, +2,20% +2,20 w isNow For basis of Wa let X = (x,y,z) EW then xtz=0 2=-Z. sol d = (A2,4,2) (0,4,0) + (-2,0,2) =y(0,1,0) + (-1,0,1). Basts of w=